OpenStudy (anonymous):

The band has 30 more members than the school chorale. If each group had 10 more members, the ratio of their membership would be 3:2. How many members are in the band?

OpenStudy (anonymous):

b = c + 30 b + 10 = 3(c + 10)/2 c + 40 = 3(c + 10)/2 2c + 80 = 3c + 30 50 = c b = 80, c = 50

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

no problem.